# Enhancing Training Efficiency Using Packing with Flash Attention

### v4

@@ Abstract
Padding is often used in tuning LLM models by adding special tokens to shorter training examples to match the length of the longest sequence in each batch. While this ensures uniformity for batch processing, it introduces inefficiencies by including irrelevant padding tokens in the computation and wastes GPU resources. Hugging Face SFT trainer has always offered the option to use packing to combine multiple training examples, allowing for maximal utilization of GPU resources. However, up till now, it did not offer proper masking of each packed training example. This capability has now been added to Hugging Face Transformers 4.43. We analyse this new feature and show the benefits across different variations of packing.
@@ 1 Introduction
To illustrate this challenge, consider a scenario where we aim to fine-tune a LLM on a corpus of text. Each sentence in the corpus may have a different length, ranging from a few words to several dozen or even hundreds. Traditional approaches would require padding shorter sequences with special tokens up to the maximum sequence length. While this ensures uniformity for processing, it introduces inefficiencies by including irrelevant padding tokens in the computation, which not only wastes GPU resources but also can dilute the model’s learning signal.
@@ 2 Related Work
There are a few ways that sample packing can be enabled with Flash Attention 2 [9] and proper attention masking of each example. The Flash Attention repository itself offers a way to pack while enabling proper masking of examples with Flash Attention. See Flash Attention Closed Issue 654 [1]. Others have proposed padding-free transformers, such as [6, 7]. The padding-free transformer methods require substantial and intrusive changes however to Hugging Face transformers library. As such, these methods have seen less uptake by the community and are not currently available in the Hugging Face library. For Hugging Face SFT Trainer users, it is desirable to have a readily available solution without requiring going outside the library. This feature is now available in Transformers version 4.43.
@@
We mention also the LengthGroupedSampler function [4] in Hugging Face Transformers library which is often used in conjunction with sample padding. We consider this method as another baseline, referred to as GroupByLength+Padding.
@@ 3 Packing with Position IDs
3.1 Online Mini-batch Collating
In mini-batch collating, the padding-free collator must pack the examples online, for each mini-batch, into a tensor of dimension, $\dim_{mini}$:
@@ 3.1 Online Mini-batch Collating
This capability has been added to Hugging Face trainer from Transformers 4.43, leveraging the new DataCollatorWithFlattening.
@@ 3.3 Collating with Optimised Sample Selection
It is possible to leverage the bin-packing-type sample selection algorithms such as [5] in conjunction with the solution provided by Packing with PositionIDs. In this case the samples to assign to each gpu and the position IDs for those samples are grouped per gpu. We call this Multipack+Position IDs. To contrast with the Multipack bin packing method, we introduce two simple baselines: (q) RandomPacking+PosID and (b) SortedPacking+PosID; Examples are packed on a first-come-first-serve basis from randomly ordered training data set and sorted (length-wise from long to short) data set, respectively.
@@ 3.4 PaddingFreeCollator
We further provide a new off-the-shelf data collator, the PaddingFreeCollator, summarised below.
@@ 3.5 Example
The modifications required are lightweight and are limited to providing the position IDs to Flash Attention. For details, see the repository at [2]. This solution relies, however, on the model exposing position IDs. As of the time of writing, 14 models expose position IDs and are supported by the solution. Specifically: llama, mistral, mixtral, granite, dbrx, falcon, gemma, olmo, phi/phi2, phi3, qwen2, qwen2_moe, stablelm, starcoder2 are all supported by the solution.
@@ 4.1 Main Results
Four approaches were evaluated. The two base cases are: (i) No packing with truncate, i.e., padding, and (ii) Basic packing (as is currently available in SFT trainer, without position IDs). Then, we evaluate two variants of the proposed solution: (iii) Our solution providing packing with position IDs, in which the batch is prepared offline and flattened, and (iv) Our solution with online packing with position IDs for each mini-batch.
@@
Table 2 provides the main results on a 20K subset of the FLAN instruct tuning dataset [15] and Table 3 shows the analogous results on the OrcaMath dataset [13]. For both datasets, all models are run for one epoch with gradient accumulation steps (gas) of 2, maximum sequence length (msl) of 4096 and mini-batch size (bs) per GPU set to 4.
Observe that the Packing with PositionID methods improve substantially the throughput in terms of tokens per second, in many case above and beyond even basic packing that does not account for improper cross-attention (the second row in each model section).
However, this maximal packing has an impact on the loss behaviour. Due to the fact that far fewer optimisation steps are taken with such maximal packing, the loss does not decrease as fast, and its effect is confirmed by the validation loss ("VLoss") after one epoch. Hence, as a best-of-both worlds remedy, we propose the online minibatch approach to Packing with PositionIDs, which due to the packing of fewer examples in each pack, does not achieve the maximal throughput, however it achieves the same optimal loss pattern and hence validation loss, as the inefficient padding-based approach. It also improves on throughput substantially beyond the padding-based approach.
@@
Table 2: Summary of main results on a 20K sample of the FLAN dataset over 10 of the 14 supported models. The table compares padding each example (first row, each section) with basic packing without position IDs (second row, each section), and two variants of the solution with position IDs: offline flattening and the online mini-batch.
@@ 4.2 Ablation Study on Minibatch Size
Figure 2: Minibatch packing offers lower GPU peak memory usage than padding approach as mini-batch size increases. “OOM" denotes out of memory error. Lower is better.
Figure 3: Minibatch packing offers better throughput (Tokens/s) than padding as mini-batch size increases. “OOM" denotes out of memory error. Higher is better.
@@ 5 Conclusions
We provide a lightweight solution for packing with PositionIDs that has been integrated into Hugging Face Transformers library version 4.43 and we demonstrate its benefits.

### v5

@@ Abstract
Padding is often used in tuning LLM models by adding special tokens to shorter training examples to match the length of the longest sequence in each batch. While this ensures uniformity for batch processing, it introduces inefficiencies by including irrelevant padding tokens in the computation and wastes GPU resources. Hugging Face trainer has always offered the option to use packing to combine multiple training examples, allowing for maximal utilization of GPU resources. However, up till now, it did not offer proper masking of each packed training example. This capability has now been added to Hugging Face Transformers 4.44. We analyse this new feature and show the benefits across different variations of packing.
@@ 1 Introduction
To illustrate this challenge, consider a scenario where we aim to fine-tune an LLM on a corpus of text. Each sentence in the corpus may have a different length, ranging from a few words to several dozen or even hundreds. Traditional approaches would require padding shorter sequences with special tokens up to the maximum sequence length. While this ensures uniformity for processing, it introduces inefficiencies by including irrelevant padding tokens in the computation, which not only wastes GPU resources but also can dilute the model’s learning signal.
@@ 2 Related Work
There are a few ways that sample packing can be enabled with Flash Attention 2 [9] and proper attention masking of each example. The Flash Attention repository itself offers a way to pack while enabling proper masking of examples with Flash Attention. See Flash Attention Closed Issue 654 [1]. Others have proposed padding-free transformers, such as [6, 7]. The padding-free transformer methods require substantial and intrusive changes however to Hugging Face transformers library. As such, these methods have seen less uptake by the community and are not currently available in the Hugging Face library. For Hugging Face Trainer users, it is desirable to have a readily available solution without requiring going outside the library. This feature is now available in Transformers version 4.44.
@@
A related effort for pre-training is studied by [pack-pretrain] who note that using causal loss masking across different documents degrades performance of the models. The authors propose to limit causal masking to within documents during pre-training. We mention also the LengthGroupedSampler function [4] in Hugging Face Transformers library which is often used in conjunction with sample padding. We consider this method as another baseline, referred to as GroupByLength+Padding.
@@ 3 Packing with Position IDs
3.1 Online minibatch Collating
In minibatch collating, the padding-free collator must pack the examples online, for each minibatch, into a tensor of dimension, $\dim_{mini}$:
@@ 3.1 Online minibatch Collating
This capability has been added to Hugging Face Trainer from Transformers 4.44, leveraging the new DataCollatorWithFlattening. Hugging Face users of the TRL library can also benefit from this capability by setting a new flag padding_free=true in the DataCollatorForCompletionOnlyLM function.
@@ 3.3 Collating with Optimised Sample Selection
One may wish to leverage bin-packing-type sample selection algorithms such as [5] in conjunction with the solution provided by Packing with PositionIDs. In this case the samples to assign to each gpu and the position IDs for those samples are grouped per gpu. We call this Multipack+Position IDs. To contrast with the Multipack bin packing method, we introduce two simple baselines: (q) RandomPacking+PosID and (b) SortedPacking+PosID; Examples are packed on a first-come-first-serve basis from randomly ordered training data set and sorted (length-wise from long to short) data set, respectively.
@@ 3.4 PaddingFreeCollator
We provide a new off-the-shelf data collator, the PaddingFreeCollator, summarised below.
@@ 3.5 Example
The modifications required are lightweight and are limited to providing the position IDs to Flash Attention. For details, see the repository at [2]. This solution relies, however, on the model exposing position IDs. As of the time of writing, 14 models expose position IDs and are supported by the solution. Specifically: llama, mistral, mixtral, granite, dbrx, falcon, gemma, olmo, phi, phi2, phi3, qwen2, qwen2_moe, stablelm, starcoder2 are all supported by the solution.
@@ 4.1 Main Results
Four approaches were evaluated. The two base cases are: (i) No packing with truncate, i.e., padding, and (ii) Basic packing (as is currently available in SFT trainer, without position IDs). Then, we evaluate two variants of the proposed solution: (iii) Our solution providing packing with position IDs, in which the batch is prepared offline and flattened, and (iv) Our solution with online packing with position IDs for each minibatch.
@@
Table 2 provides the main results on a 20K subset of the FLAN instruct tuning dataset [15] and Table 3 shows the analogous results on the OrcaMath dataset [13]. For both datasets, all models are run for one epoch with gradient accumulation steps (gas) of 2, maximum sequence length (msl) of 4096 and minibatch size (bs) per GPU set to 4.
Observe that the Packing with PositionID methods improve the throughput substantially in terms of tokens per second, in many case above and beyond even basic packing that does not account for improper cross-attention (the second row in each model section).
However, this maximal packing has an impact on the loss behaviour. Due to the fact that far fewer optimisation steps are taken with such maximal packing, the loss does not decrease as fast, and its effect is confirmed by the validation loss ("VLoss") after one epoch. Hence, as a best-of-both worlds remedy, we propose the online minibatch approach to Packing with PositionIDs. Due to the packing of fewer examples in each pack, minibatch packing does not achieve the maximal throughput, however it achieves the same optimal loss pattern and hence validation loss, as the inefficient padding-based approach. It also improves on throughput substantially beyond the padding-based approach.
@@
Table 2: Summary of main results on a 20K sample of the FLAN dataset over 10 of the 14 supported models. The table compares padding each example (first row, each section) with basic packing without position IDs (second row, each section), and two variants of the solution with position IDs: offline flattening and the online minibatch.
@@ 4.2 Ablation Study on Minibatch Size
Figure 2: Minibatch packing offers lower GPU peak memory usage than padding approach as minibatch size increases. “OOM" denotes out of memory error. Lower is better.
Figure 3: Minibatch packing offers better throughput (Tokens/s) than padding as minibatch size increases. “OOM" denotes out of memory error. Higher is better.
@@ 5 Conclusions
We provide a lightweight solution for packing with PositionIDs that has been integrated into Hugging Face Transformers library version 4.44 and we demonstrate its benefits.